OpenStudy (akohl):

@Abhisar A spring-loaded dart gun is compressed a distance d, then fired straight up into the air. The dart reaches a maximum height of 24 m. For a second trial, the spring is compressed a distance of 0.50d. What is the maximum height of the dart in the second trial? Assume that the effect of friction and other nonconservative forces is negligible. 3.0 m 12 m 6.0 m 48 m

OpenStudy (akohl):

I think its 12 not sure

OpenStudy (abhisar):

What happened about the yesterday's question ?

OpenStudy (akohl):

Teacher said it was 554 that's why it was relative I the air not ground

OpenStudy (akohl):

To the air

OpenStudy (abhisar):

That means my answer was correct ?

OpenStudy (akohl):

Yep thanks

OpenStudy (abhisar):

Alas !

OpenStudy (akohl):

Is it 12 because its compressed to half the length

OpenStudy (abhisar):

one min !

OpenStudy (abhisar):

Yes that's correct. But let me show you a mathematical way to calculate/prove it.

OpenStudy (abhisar):

The potential energy stored in the spring while we compress it is transferred to the dart which gets converted into its kinetic energy. The maximum height attained by the dart will only be due to this energy. Now in \(\huge\sf Case~1\) \(\sf \huge\frac{1}{2}k.d^2=mg\times24\\~\\~\\\large k=48mg/d^2\)

OpenStudy (abhisar):

Now, in case 2 \(\large\sf \frac{1}{2}.k.(0.5d)^2=mgh\\~\\~substitute~the~value~of~k~and~find~h\)

OpenStudy (akohl):

I got 12 thanks

OpenStudy (abhisar):

\(\color{red}{\huge\bigstar}\huge\text{You're Most Welcome! }\color{red}\bigstar\) \(~~~~~~~~~~~~~~~~~~~~~~~~~~~\color{green}{\huge\ddot\smile}\color{blue}{\huge\ddot\smile}\color{pink}{\huge\ddot\smile}\color{red}{\huge\ddot\smile}\color{yellow}{\huge\ddot\smile}\)

OpenStudy (abhisar):

Maximum height will be 6m. This means you didn't solved it !

OpenStudy (abhisar):

\(\color{blue}{\text{Originally Posted by}}\) @Abhisar Now, in case 2 \(\large\sf \frac{1}{2}.k.(0.5d)^2=mgh\\~\\~substitute~the~value~of~k~and~find~h\) \(\color{blue}{\text{End of Quote}}\) \(\large \sf \frac{1}{2}.\frac{48mg}{d^2}\times\frac{d^2}{4}=mgh\) So h=6m @akohl